OpenStudy (anonymous):

Triangle GHK is similar to triangle SDW and the ratio of each pair of corresponding sides of the triangles is 3:4. If the length of altitude HI is x-3 and the length of altitude DJ is x-1, find the lengths of HI and DJ.

OpenStudy (anonymous):



OpenStudy (anonymous):

I think you have made a mistake. is not that triangly GHK is 4 and triangle SDW is 3?

OpenStudy (anonymous):

If that is the case than in a more simplified ratio is 1:0.75, to say that\[x+3=0.75(x-1)\]

OpenStudy (anonymous):

x+3=0.75x-0.75 0.25x=-3.75 x=-15

OpenStudy (anonymous):

3:4 is the ratio of the corresponding pairs